Emile wants to win the 100 yard dash at the track meet. If he runs 4.9 miles per hour, how many minutes will it take him to run the race?
Answer by HyperBrain(694)

4.9 mi/hr= (4.9)(1760 yd)/(60 minutes)=2156/15 yd/min.
---
time=distance/speed=100 yd/ (2156/15 yd/min)= 0.6957 min